OpenStudy (clara1223):

List the x-intercepts for y=cos((1/2)x+(pi/5)) on the interval [(-2pi/5), (4pi)]

OpenStudy (clara1223):

\[y=\cos(\frac{ 1 }{ 2 }x+\frac{ \pi }{ 5 })\] on the interval \[\left[ \frac{ -2\pi }{ 5 }, 4\pi \right]\]